# CURSOR: Scalable Mixed-Order Hypergraph Matching with CUR Decomposition

### 26 Feb 2024 v1

@@ 1 Introduction
Finding the correspondences between a pair of feature sets by graph-matching has many applications in computer vision and pattern recognition tasks like feature tracking [8, 32], image classification [29], object detection [19], and gene-drug association identification [3]. The second-order graph matching (pairwise matching) problem is a quadratic assignment problem (QAP), which is NP-hard [14]. Efforts to find soft-constraint approximate solutions [17, 6, 5, 25] have been limited to the representation of pairwise compatibilities.
Higher-order graphs, known as hypergraphs [7, 16, 15, 24, 9], integrate better geometric information and handle transformations like scaling and rotation better. The hypergraph matching problem considers the compatibility between two hypergraphs as a high-order tensor. The objective function aims to find the maximization over all permutations of the features. A kd-tree-based approximate nearest neighbor (ANN) method [7] to reduce the space and time complexity of the hypergraph matching algorithm has been used in many hypergraph matching algorithms (e.g. RRWHM [16], BCAGM [24] and ADGM [15]). ANN-based methods compute the sparse compatibility tensor by searching for nearest neighbors between randomly sampled hyperedges in the source graph and all the hyperedges in the target graph. Large-scale $k^{\text{th}}$-order hypergraph matching with $n_{2}$ points in the target graph has $O(n_{2}^{k})$ time and space complexity when calculating the compatibilities for each source graph hyperedge. Achieving a higher matching accuracy requires as many of the highest compatibilities as possible to increase the probability of finding the ground truth paired hyperedges, resulting in a denser compatibility tensor.
@@ 2.1 Hypergraph Matching
Earlier second-order graph-matching algorithms represented the geometric consistency between a pair of features as the edges of a graph to avoid ambiguities like repeated patterns and textures. These algorithms pursued an approximate solution as the problem is known to be NP-hard. Among them, Leordeanu and Herbert [17] and Cour et al. [6] estimated the rank-1 approximation of the compatibility matrix with power iteration as the flattening of the assignment matrix. Cho et al. [5] provided a novel random walk view for graph matching with a reweighting jump scheme and reduced the constraints in its iterative process. Recently, Wang et al. [25] proposed a functional representation for graph matching with the additional goal of avoiding the compatibility matrix construction. The performance of second-order methods was limited with the restriction to the normal graph embedding pairwise relationships.
In the past decade, to overcome the limitation of pairwise similarity, various hypergraph matching algorithms were developed based on the structural compatibilities between the higher-order hyperedges of two hypergraphs. In order to avoid the enormous computational cost of a full compatibility tensor, Duchenne et al. [7], extending the SM algorithm proposed by [17] to a higher order, constructed the sparse compatibility tensor with an ANN-based method, which is frequently used for generating compatibility tensors in later hypergraph matching works. Lee et al. [16] introduced the method of [5] to reweighted walk hypergraph matching problems, enforcing the matching constraint with a bi-stochastic normalization scheme. Lê-Huu and Paragios [15] decomposed the problem under different constraints and handled it with an alternating direction method of multipliers. Khan et al. [12] directly applied CUR decomposition to the full compatibility matrix and tensor at the cost of a higher space complexity than the ANN-based method. Although the ANN-based method can significantly decrease the time complexity during the matching process, it still has an enormous computational cost to calculate the compatibility tensor.
With the rise of deep learning, learning-based graph-matching algorithms were proposed to learn the parameters as deep feature extraction hierarchies in a data-driven way [31, 27, 26, 20]. Inspired by Wang et al. [26], Liao et al. proposed the first unified hypergraph neural network, HNN-HM. Unlike the sparse compatibility tensor widely used in ANN-based algorithms, the dense deep feature matrices or tensors require more computation resources. Therefore, compared with ANN-based algorithms, the learning-based hypergraph matching algorithms can only handle much smaller-scale problems under the same memory constraint.
@@ 3.3 Fiber-CUR-based Tensor Generation
The ANN-based method searches for the highest compatibilities in all target hyperedges, which is time and space-consuming for large-scale hypergraph matching. Based on randomized fiber CUR decomposition, the third-order compatibility tensor can be generated with the pairwise graph-matching result to reduce the computational cost. The second-order matching result from Algorithm 1 is $\mathcal{P}^{k}$, where $k\ll n_{2}$. The method does not need to compare the hyperedge $(i_{1},i_{2},i_{3})\in\mathcal{E}_{1}$ with all the hyperedges in the target hypergraph, but with the hyperedge set $\{(j_{1},j_{2},:),(j_{1},:,j_{3}),(:,j_{2},j_{3})\}\in\mathcal{E}_{2}$ where $j_{1}\in\mathcal{P}_{i_{1}}^{k}$, $j_{2}\in\mathcal{P}_{i_{2}}^{k}$, and $j_{3}\in\mathcal{P}_{i_{3}}^{k}$. The $r$ highest entries from each hyperedge’s calculated compatibilities in the source graph are selected. Since each hyperedge is compared with fewer target hyperedges, $r$ is much smaller than the number of selected compatibilities $r_{1}$ in [7].
@@ 4.1 Large-Scale Random Synthetic Dataset
CURSOR was evaluated on synthetic datasets with increasing problem scales. Results of the ADGM algorithm based on ANN, whose accuracy was the highest among all the prior state-of-the-art hypergraph matching algorithms, were given. The parameter settings of ADGM, including $t$ and $r_{1}$, strictly followed the original work [15]. CURSOR can deal with large-scale scenarios and achieve privilege-matching accuracy with much less memory usage (Table 1). CURSOR was capable of solving 1000-vs-1000 matching problems with high accuracy, while ADGM failed to generate the tensor when $n_{1}>100$ under the same memory constraint.
@@ 4.3 CMU House and Hotel Dataset
The sequence-matching results are given in Fig. 3. All the learning-free-based matching algorithms combined with CURSOR surpassed the original ANN-based ones except for the BCAGM3 algorithm. The BCAGM3 algorithm with CURSOR frequently made descent errors during the experiment as its matching accuracy heavily decreased when the compatibility tensor was too sparse. CURSOR obtained comparable matching accuracy to the state-of-the-art learning-free algorithms when the outlier number was ten and achieved the best matching accuracy under the 30-vs-30 case on the House dataset. For the Hotel dataset, CURSOR achieved a much higher matching accuracy than ANN, but was less accurate than HNN-HM on 20-vs-30 problems as HNN-HM was specifically trained on the House Dataset. It is noteworthy that CURSOR solved the matching problem with a unified model while HNN-HM must first train separating models in different datasets to achieve a better matching result.
@@ 4.4 Car and Motorbike Dataset
The Car and Motorbikes dataset introduced in [18] consists of 30 real-life car image pairs and 20 motorbike image pairs, and was used in previous works to evaluate matching algorithms [16, 15, 24]. In this experiment, all inlier points in both images were kept and labeled outlier points were randomly chosen in the second image, with the number varying from 0 to 30 in stpdf of 5. Every image pair in both datasets was matched. $t=n_{1}n_{2}$ hyperedges were selected in the first image. For ANN-based methods, $r_{1}$ was set as $0.3n_{1}n_{2}$. CURSOR selected $r=50$ highest compatibilities in each tensor block. During experiments $k=10$ and $c=300$
Figure 3(a) reported the average number of non-zero elements in the sparse compatibility tensors generated by the two types of methods. The average accuracy with the ANN-based methods and CURSOR was shown in Fig. 3(b) and Fig. 3(c). All algorithms integrated with CURSOR consistently improved their matching performance with a compatibility tensor more than ten times sparser than the ANN-based methods. In most cases, the default PRL-based matching algorithm achieved higher accuracy than other state-of-the-art algorithms. Figure 5 shows matching examples. When combined with CURSOR, other hypergraph matching algorithms showcased fewer mismatches. CURSOR, with the PRL-based algorithm, obtained the highest matching accuracy.
@@ 5 Conclusion
We propose CURSOR, a cascaded mixed-order hypergraph matching framework based on CUR decomposition for scalable graph matching. The framework contains a CUR-based second-order graph matching algorithm and a fiber-CUR-based tensor generation method, which significantly decreases the computational cost, and can be seamlessly integrated into existing state-of-the-art hypergraph matching algorithms to enhance their performance. A PRL-based hypergraph matching algorithm for sparse compatibility tensors is developed to accelerate the convergence. Experiment results demonstrated that CURSOR contributes to a higher matching accuracy with a sparser compatibility tensor, which has more potential utility in the big-data era. In future work, we plan to develop a more principled adaptive scheme to optimize the parameters of CURSOR so that it can perform better with fewer computations on larger-scale tasks.
@@ 6 PRL-based Method Derivation
Assume the highest probability for $i_{0}\in\mathcal{V}_{1}$ is $p_{i_{0}}(j_{0})$. [10] proved that $p_{i_{0}}^{(k)}(j_{0})$ consistently satisfies $\sum_{l=1}^{n_{1}}\sum_{m=1}^{n_{2}}R_{i_{0}l}(j_{0},m)p_{l}^{(k)}(m)\geq p_{i_{0}}^{(k)}(j_{0})$ during the updating if $\mathbf{R}$, the matrix form of $R_{i_{1}i_{2}}(j_{1},j_{2})$ for all $(i_{1},i_{2})$ and $(j_{1},j_{2})$, is symmetric. To further accelerate the convergence, the weighting factor is replaced from $p_{i}^{(k)}(j)$ to $\sum_{l=1}^{n_{1}}\sum_{m=1}^{n_{2}}R_{il}(j,m)p_{l}^{(k)}(m)$, which leads to
@@
where the first-order compatibility vector $\hat{\mathbf{m}}$ is obtained by column-wise flattening $\hat{\mathbf{M}}$ and $\mathbf{H}$ is the second-order compatibility matrix. Since both $\mathbf{H}$ and $\hat{\mathbf{m}}$ are dense, with the all-ones vector as $\mathbf{x}^{(0)}$, Eq. (19) converges consistently.
@@
where $\odot$ is the element-wise multiplication and $\alpha\in[0,1]$ is a balance weight between the first and third-order compatibilities. $\otimes_{l}$ is the mode-$l$ product of the tensor and vector. Since the third-order compatibility tensor $\mathcal{H}$ is highly sparse, the high value of $\hat{\mathbf{x}}^{(k+1)}$ in (20) is concentrated if the sparse tensor is reliable. In our work, a reliable tensor means most ground truth hyperedge pair compatibilities are successfully selected in the tensor blocks.
7 Ablation Studies
7.1 CUR-based Pairwise Matching
Figure 6: 5(a) CUR-based second-order graph matching accuracy with various $c$. 5(b) The average accuracy using CURSOR with CUR-based pairwise matching result and CURSOR with randomly sampled indices (CURSOR-R).
Method Time (s)
Computing $\mathbf{H}$ Matching
Full Matrix 6.311 0.491
$c=10$ 0.018 0.016
$c=30$ 0.031 0.016
$c=100$ 0.109 0.043
The CUR-based second-order graph matching was first evaluated on the dataset with various $c$, i.e., the number of randomly selected columns from $\mathbf{H}$. During the experiment, only the second-order graph-matching result is analyzed. We compared the proposed method with a PRL-based pairwise matching using the full compatibility matrix. The matching accuracy is shown in Fig. 5(a). Due to the low-rank estimation, the CUR-based method decreased the matching performance. Specifically, as the columns were randomly chosen, the result was quite poor when $c$ was relatively small. However, with less than $1\%$ (100 out of 100$\times$110) of columns selected, the CUR-based pairwise matching algorithm achieved a comparable result to the algorithm using the whole matrix. Table 2 reports the average time consumption of the compatibility matrix generation and graph matching. With only a few columns calculated, the CUR-based second-order algorithm effectively accelerated the matching process.
The effectiveness of the intermediate second-order matching result was further studied. As discussed in the main paper, the second-order matching result $\mathcal{P}^{k}=\{\mathcal{P}_{1}^{k},\cdots,\mathcal{P}_{n_{1}}^{k}\}$ consists $k$ best-matching target nodes for each source node, where $n_{1}$ denotes the number of the source nodes. The hypergraph matching result with $\mathcal{P}^{k}$ was compared to the result with $k_{1}$ randomly sampled indices in all three tensor modes (denoted as CURSOR-R). The parameter $c$ and $k$ from $\mathcal{P}^{k}$ for CURSOR with the pairwise matching result was set as 100 and 10 respectively. For CURSOR-R, $k_{1}$ varied from 10 to 50. For both methods, the number of randomly selected hyperedges from the source hyperedges $t=3000$, the number of highest compatibilities in each tensor block $r=100$, and the balance factor of PRL-based algorithm $\alpha$ was set to $0.2$. As shown in Fig. 5(b), CURSOR-R had a lower matching accuracy even for $k_{1}=50$. Assuming the number of nodes is $n_{2}$ in the target graph. Theoretically, only 3 out of $3n_{2}^{2}$ fibers for each tensor block, i.e., key fibers in our work, contain the entry of the ground truth hyperedge pair. Due to the random sampling, the probability of selecting the key fiber is $(k_{1}/n_{2})^{2}$, around $20.7\%$ when $k_{1}=50$ and $n_{2}=110$. Therefore, the sparse compatibility tensor generated by CURSOR-R was highly unreliable. CURSOR with the CUR-based pairwise matching result selected the key fiber in each tensor block with a high probability, effectively increasing the final accuracy.
7.2 CURSOR vs ANN-based Tensor Generation
Figure 7: Results on 100-vs-110 synthetic dataset comparing CURSOR with ANN, $\alpha=0.2$ for CURSOR. 6(a) The average matching accuracy using CURSOR and ANN with different $\alpha$. 6(b) The average iterations for convergence using CURSOR and ANN with different $\alpha$. 6(c) The decay $\|\mathbf{x}^{(k+1)}-\mathbf{x}^{(k)}\|_{2}$ per iteration with different $\alpha$ using CURSOR and ANN. $\sigma$ is set as 0.1.
The average matching accuracy is reported in Fig. 6(a). The ANN method shows an unstable performance with high $\sigma$ since the compatibility tensor was too sparse. When $\alpha$ was low, since the PRL-based algorithm focused on the third-order compatibilities, the ANN method did not generate a reliable compatibility tensor. To make the algorithm focus more on first-order compatibilities, $\alpha$ was further increased to 0.8 for the ANN case, which significantly improved the accuracy. The average number of iterations to converge is shown in Fig. 6(b). When $\sigma>0.02$, the ANN method did not converge within 100 iterations. With the same tensor sparsity, CURSOR successfully converged in all the cases. The decay of $\mathbf{x}^{(k)}$ per iteration for the ANN method and CURSOR was further analyzed, as shown in Fig. 6(c). The ANN method did not converge due to few ground truth hyperedge compatibilities selected from the whole tensor. However, CURSOR chose the non-zero compatibilities from a smaller reliable searching region. Therefore, it is capable of converging fast with the same tensor sparsity.
8 Parameter Sensitivity Analysis
8.1 Parameter in Second-order Graph Matching
Figure 8: 7(a) The average hit rate over increasing $k$ on a 100-vs-110 synthetic dataset with $\sigma=0.02$ from the CUR-based second-order graph matching results. 7(b) The average accuracy on the 100-vs-110 synthetic dataset with $\sigma\in[0,0.2]$ and $c\in[10,100]$.
8.2 Parameters During Tensor Generation
Figure 9: Results on 100-vs-110 synthetic dataset with $\sigma\in[0,0.2]$ using CURSOR. 8(a) The matching accuracy with increasing $k$. 8(b) The matching accuracy with increasing $r$.
Figure 10: Results on 100-vs-110 synthetic dataset with PRL-based matching algorithm. 9(a) The average matching accuracy using CURSOR with different $\alpha$. 9(b) The average iterations for convergence using CURSOR with different $\alpha$. 9(c) The decay $\|\mathbf{x}^{(k+1)}-\mathbf{x}^{(k)}\|_{2}$ per iteration with different $\alpha$ settings using CURSOR. $\sigma$ is set as 0.1.
8.3 Parameters of PRL-based Matching Algorithm
To study the sensitivity of $\alpha$, the matching accuracy and convergence speed using CURSOR were further analyzed. During the experiment, $c=100$, $k=10$, and $r$ was set as $100$. The stopping criteria of the PRL-based algorithm was assigned as $\|\mathbf{x}^{(k+1)}-\mathbf{x}^{(k)}\|_{2}\leq 10^{-8}$. The matching accuracy with various $\alpha$ is shown in Fig. 9(a). With a reliable compatibility tensor, the PRL-based matching algorithm achieved almost the same performance regardless of $\alpha$. We further analyze their convergence speeds, as shown in Fig. 9(b) and Fig. 9(c). As the noise level increased, more iterations were required to satisfy the stopping criteria, and a lower $\alpha$ achieved faster convergence. As discussed in the previous sections, the parameter $\alpha$ is a balanced factor between first and third-order compatibilities. Although the method converged faster with $\alpha=0$, the first-order compatibilities stabilized the matching process and increased the matching performance under some extreme circumstances.
9 Memory Footprint Analysis

### 14 Mar 2024 v2

@@ 1 Introduction
Finding the correspondences between a pair of feature sets by graph-matching has many applications in computer vision and pattern recognition tasks like feature tracking [11, 39, 23], image classification [36], object detection [24], and gene-drug association identification [5]. The second-order graph matching (pairwise matching) problem is a quadratic assignment problem (QAP), which is NP-hard [18]. Efforts to find soft-constraint approximate solutions [21, 8, 7, 32] have been limited to the representation of pairwise compatibilities.
Higher-order graphs, known as hypergraphs [9, 20, 19, 30, 12], integrate better geometric information and handle transformations like scaling and rotation better. The hypergraph matching problem considers the compatibility between two hypergraphs as a high-order tensor. The objective function aims to find the maximization over all permutations of the features. A kd-tree-based approximate nearest neighbor (ANN) method [9] to reduce the space and time complexity of the hypergraph matching algorithm has been used in many hypergraph matching algorithms (e.g. RRWHM [20], BCAGM [30] and ADGM [19]). ANN-based methods compute the sparse compatibility tensor by searching for nearest neighbors between randomly sampled hyperedges in the source graph and all the hyperedges in the target graph. Large-scale $k^{\text{th}}$-order hypergraph matching with $n_{2}$ points in the target graph has $O(n_{2}^{k})$ time and space complexity for the compatibilities of each source graph hyperedge. Achieving a higher matching accuracy requires as many of the highest compatibilities as possible to increase the probability of finding the ground truth paired hyperedges, resulting in a denser compatibility tensor.
@@ 2.1 Hypergraph Matching
Second-order graph-matching algorithms [21, 8, 7, 32, 10] represented the geometric consistency between a pair of features as the edges of a graph to avoid ambiguities like repeated patterns and textures. These algorithms pursued an approximate solution as the problem is known to be NP-hard. Among them, Leordeanu and Herbert [21] and Cour et al. [8] estimated the rank-1 approximation of the compatibility matrix with power iteration as the flattening of the assignment matrix. Cho et al. [7] provided a novel random walk view for graph matching with a reweighting jump scheme and reduced the constraints in its iterative process. Recently, Wang et al. [32] proposed a functional representation for graph matching with the additional goal of avoiding the compatibility matrix construction. The performance of second-order methods was limited with the restriction to the normal graph embedding pairwise relationships.
In the past decade, to overcome the limitation of pairwise similarity, various hypergraph matching algorithms [9, 20, 19, 12] were developed based on the structural compatibilities between the higher-order hyperedges of two hypergraphs. In order to avoid the enormous computational cost of a full compatibility tensor, Duchenne et al. [9], extending the SM algorithm proposed by [21] to a higher order, constructed the sparse compatibility tensor with an ANN-based method, which is frequently used for generating compatibility tensors in later hypergraph matching works. Lee et al. [20] introduced the method of [7] to reweighted walk hypergraph matching problems, enforcing the matching constraint with a bi-stochastic normalization scheme. Lê-Huu and Paragios [19] decomposed the problem under different constraints and handled it with an alternating direction method of multipliers. Khan et al. [16] directly applied CUR decomposition to the full compatibility matrix and tensor at the cost of a higher space complexity than the ANN-based method. Although the ANN-based method can significantly decrease the time complexity during the matching process, it still has an enormous computational cost to calculate the compatibility tensor.
With the rise of deep learning, various learning-based graph-matching algorithms were proposed to learn the parameters as deep feature extraction hierarchies in a data-driven way [38, 34, 26, 3, 31, 15, 4]. Inspired by Wang et al. [33], Liao et al. [25] proposed the first unified hypergraph neural network, HNN-HM. Unlike the sparse compatibility tensor widely used in ANN-based algorithms, the dense deep feature matrices or tensors require more computation resources. Therefore, compared with ANN-based algorithms, the learning-based hypergraph matching algorithms can only handle much smaller-scale problems under the same memory constraint.
@@ 3.3 Fiber-CUR-based Tensor Generation
The ANN-based method searches for the highest compatibilities in all target hyperedges, which is time and space-consuming for large-scale hypergraph matching. Based on randomized fiber CUR decomposition, the third-order compatibility tensor can be generated with the pairwise graph-matching result to reduce the computational cost. The second-order matching result from Algorithm 1 is $\mathcal{P}^{k}$, where $k\ll n_{2}$. The method does not need to compare the hyperedge $(i_{1},i_{2},i_{3})\in\mathcal{E}_{1}$ with all the hyperedges in the target hypergraph, but with the hyperedge set $\{(j_{1},j_{2},:),(j_{1},:,j_{3}),(:,j_{2},j_{3})\}\in\mathcal{E}_{2}$ where $j_{1}\in\mathcal{P}_{i_{1}}^{k}$, $j_{2}\in\mathcal{P}_{i_{2}}^{k}$, and $j_{3}\in\mathcal{P}_{i_{3}}^{k}$. The $r$ highest entries from each hyperedge’s calculated compatibilities in the source graph are sellected. Since each hyperedge is compared with fewer target hyperedges, $r$ is much smaller than the number of selected compatibilities $r_{1}$ in [9].
@@ 4.1 Large-Scale Random Synthetic Dataset
CURSOR was evaluated on synthetic datasets with increasing problem scales. Results of the ADGM algorithm based on ANN, whose accuracy was the highest among all the prior state-of-the-art hypergraph matching algorithms, were given. The parameter settings of ADGM, including $t$ and $r_{1}$, strictly followed the original work [19]. CURSOR can deal with large-scale scenarios and achieve privilege-matching accuracy with much less memory usage (Table 1). CURSOR was capable of solving 1000-vs-1000 matching problems with high accuracy, while ADGM failed to generate the tensor when $n_{1}>100$ under the same memory constraint. A more detailed memory footprint analysis and the potential bottleneck of CURSOR will be provided in the supplement.
@@ 4.3 CMU House and Hotel Dataset
The sequence-matching results are given in Fig. 3. All the learning-free-based matching algorithms combined with CURSOR surpassed the original ANN-based ones except for the BCAGM3 algorithm. The BCAGM3 algorithm with CURSOR frequently made descent errors during the experiment as its matching accuracy heavily decreased when the compatibility tensor was too sparse. CURSOR obtained comparable matching accuracy to the state-of-the-art learning-free algorithms when the outlier number was ten and achieved the best matching accuracy under the 30-vs-30 case on the House dataset. For the Hotel dataset, CURSOR achieved a much higher matching accuracy than ANN, but was less accurate than HNN-HM on 20-vs-30 problems as HNN-HM was specifically trained on the House Dataset. It is noteworthy that CURSOR solved the matching problem with a unified model while HNN-HM must first train separating models in different datasets to achieve a better matching result. More detailed analysis will be provided in the supplement to show the effectiveness of CURSOR.
@@ 4.4 Car and Motorbike Dataset
The Car and Motorbikes dataset [22] consists of 30 real-life car image pairs and 20 motorbike image pairs, and was used in previous works to evaluate matching algorithms [20, 19, 30]. In this experiment, all inlier points in both images were kept and labeled outlier points were randomly chosen in the second image, with the number varying from 0 to 30 in step of 5. Every image pair in both datasets was matched. $t=n_{1}n_{2}$ hyperedges were selected in the first image. For ANN-based methods, $r_{1}$ was set as $0.3n_{1}n_{2}$. CURSOR selected $r=50$ highest compatibilities in each tensor block. During experiments $k=10$ and $c=300$
Figure 3(a) reported the average number of non-zero elements in the compatibility tensors generated by the two types of methods. The average accuracy with the ANN-based methods and CURSOR was shown in Fig. 3(b) and Fig. 3(c). All algorithms integrated with CURSOR consistently improved their matching performance with a compatibility tensor more than ten times sparser than the ANN-based methods. In most cases, the default PRL-based matching algorithm achieved higher accuracy. Figure 5 shows matching examples. Combined with CURSOR, other hypergraph matching algorithms showcased fewer mismatches. CURSOR, with the PRL-based algorithm, obtained the highest matching accuracy.
@@ 5 Conclusion
We propose CURSOR, a cascaded mixed-order hypergraph matching framework based on CUR decomposition for scalable graph matching. The framework contains a CUR-based second-order graph matching algorithm and a fiber-CUR-based tensor generation method, which significantly decreases the computational cost, and can be seamlessly integrated into existing state-of-the-art hypergraph matching algorithms to enhance their performance. A PRL-based hypergraph matching algorithm for sparse compatibility tensors is developed to accelerate the convergence. Experiment results demonstrated that CURSOR contributes to a higher matching accuracy with a sparser tensor, which has more potential utility in the big-data era. In future work, we plan to develop a more principled adaptive scheme to optimize the parameters of CURSOR so that it can perform better with fewer computations on larger-scale tasks.
Acknowledgment
[…]
@@ 6 PRL-based Method Derivation
Assume the highest probability for $i_{0}\in\mathcal{V}_{1}$ is $p_{i_{0}}(j_{0})$. Hummel and Zucker proved that $p_{i_{0}}^{(k)}(j_{0})$ consistently satisfies $\sum_{l=1}^{n_{1}}\sum_{m=1}^{n_{2}}R_{i_{0}l}(j_{0},m)p_{l}^{(k)}(m)\geq p_{i_{0}}^{(k)}(j_{0})$ during the updating if $\mathbf{R}$, the matrix form of $R_{i_{1}i_{2}}(j_{1},j_{2})$ for all $(i_{1},i_{2})$ and $(j_{1},j_{2})$, is symmetric [13]. To further accelerate the convergence, the weighting factor is replaced from $p_{i}^{(k)}(j)$ to $\sum_{l=1}^{n_{1}}\sum_{m=1}^{n_{2}}R_{il}(j,m)p_{l}^{(k)}(m)$, which leads to
@@
where $\odot$ is the element-wise multiplication and the first-order compatibility vector $\hat{\mathbf{m}}$ is obtained by column-wise flattening $\hat{\mathbf{M}}$. The square calculation in Eq. (19) is also element-wise. $\mathbf{H}$ is the second-order compatibility matrix. Since both $\mathbf{H}$ and $\hat{\mathbf{m}}$ are dense, with the all-ones vector as $\mathbf{x}^{(0)}$, Eq. (19) converges consistently.
@@
where $\otimes_{l}$ is the mode-$l$ product of the tensor and vector and $\alpha\in[0,1]$ is a balance weight between the first and third-order compatibilities. Since the third-order compatibility tensor $\mathcal{H}$ is highly sparse, the high value of $\hat{\mathbf{x}}^{(k+1)}$ in Eq. (20) is concentrated if the sparse tensor is reliable. In our work, a reliable tensor means most ground truth hyperedge pair compatibilities are successfully selected in the tensor blocks.
7 Detailed Analysis in Sec. 4
Due to the space limitation of the main paper, we provide a more detailed experiment analysis based on the results of Sec. 4 to discuss the superiority and bottleneck of CURSOR.
7.1 Memory Footprint Analysis in Sec. 4.1